# One-cycle Structured Pruning with Stability Driven Structure Search

### v1

@@ Abstract
Existing structured pruning typically involves multi-stage training procedures that often demand heavy computation. Pruning at initialization, which aims to address this limitation, reduces training costs but struggles with performance. To address these challenges, we propose an efficient framework for one-cycle structured pruning without compromising model performance. In this approach, we integrate pre-training, pruning, and fine-tuning into a single training cycle, referred to as the ‘one cycle approach’. The core idea is to search for the optimal sub-network during the early stages of network training, guided by norm-based group saliency criteria and structured sparsity regularization. We introduce a novel pruning indicator that determines the stable pruning epoch by assessing the similarity between evolving pruning sub-networks across consecutive training epochs. Also, group sparsity regularization helps to accelerate the pruning process and results in speeding up the entire process. Extensive experiments on datasets, including CIFAR-10/100, and ImageNet, using VGGNet, ResNet, MobileNet, and ViT architectures, demonstrate that our method achieves state-of-the-art accuracy while being one of the most efficient pruning frameworks in terms of training time. The source code will be made publicly available.
@@ 1 Introduction
Over the past decade, deep neural networks (DNNs) have demonstrated superior performance at the expense of substantial memory usage, computing power, and energy consumption [21, 8, 63, 17]. For compressing and deploying these complex models on low-capability devices, pruning has emerged as a particularly promising approach for many embedded and general-purpose computing platforms. Structured pruning [32, 56, 5, 52, 13, 6] and unstructured pruning [19, 14, 11, 51] represent the two predominant methods in pruning. Also, there has been increasing interest in middle-level pruning granularity [38, 65, 47], which provides the capability to achieve fine-grained structured sparsity combining the advantages of both unstructured fine-grained sparsity and structured coarse-grained sparsity.
Figure 1: Pruning ResNet50 on the ImageNet dataset. Note that the size of the marker represents the training speedup relative to the baseline training cost.
@@
With the emergence of large-scale DNN models [3, 46, 63], there is a growing need to reassess the conventional pruning approach of pre-training, pruning, and fine-tuning, which demands excessive resources. Conventional pruning often requires multi-stage pipelines with repeated fine-tuning procedures or offline re-training. Without efficient learning mechanisms, it is not only a waste of computational power and time but also less applicable to real-time scenarios. One-cycle training and pruning, also known as one-cycle pruning, presents itself as a viable alternative solution to this problem. Obviously, it is important to explore whether a one-cycle pruning framework can rapidly learn new tasks or adapt to new environments in various real-world applications, such as autonomous vehicles, drones, robots, and so on [15, 5].
To address the limitations of conventional pruning, we introduce One-Cycle Structured Pruner (OCSPruner), which eliminates the requirement for pre-trained base network for pruning. Considering scenarios where networks need to learn from scratch within limited resource budgets, the efficiency gained from avoiding pre-training the base model becomes more compelling. Hence, we focus on efficient and stable structured pruning, ensuring model performance in a more sustainable way. Basically, structured pruning offers the advantage of creating slimmer yet structured networks that do not rely on specific hardware or software; however, it can occasionally lead to performance drops, particularly when dealing with high pruning ratios. Considering the insights from the Lottery Ticket Hypothesis [14], which suggests that sparse, trainable sub-networks can be identified within larger networks, the timing and method of pruning become critical. Our focus on one-cycle structured pruning involves pruning the network at initialization [14, 30, 7]. However, this type of method may suffer from potential sub-optimal network structures that result in performance degradation. Thus, rather than pruning at initialization, our proposed strategy involves early-phase pruning during training while still maintaining a one-cycle approach. Despite related methods like pruning-aware training (PaT) [52] and loss-aware automatic selection of pruning criteria (LAASP) [18], they overlook structural parameter grouping during saliency estimation, leading to reduced performance [13]. Specifically, PaT [52] focuses solely on total channel numbers, not considering individual channel identities during stable sub-network selection, while LAASP [18] lacks stability estimation criteria and involves complex iterative pruning. This paper addresses these limitations and introduces additional algorithmic enhancements for early-phase structured pruning during training from scratch, resulting in state-of-the-art performance.
@@
The main contributions of this work are summarized as follows:
@@
  • A robust pruning paradigm is suggested, which continuously monitors the evolving pruning sub-structure guided by a pruning stability indicator to achieve an optimal pruning architecture during the early stages of network training.
  • The effectiveness of the proposed algorithm is validated in various datasets utilizing single-branch VGGNet, multi-branch ResNet, and compact MobileNet architectures. Notably, it achieves 75.49% top-1 and 92.63% top-5 accuracy for the ResNet50 model on the ImageNet dataset, demonstrating a 1.38×1.38$\times$1.38 × training speedup compared to baseline training while reducing network FLOPs by over 57%.
@@ 2 Related Work
Structured Pruning. While unstructured pruning has been a subject of long-standing research [28, 20, 19], structured pruning gained popularity with the emergence of modern DNNs [32]. Structured pruning revolves around extracting efficient sub-structures from complete models while optimizing storage and inference processes. This can involve incorporating regularization techniques to induce sparsity [44, 61] or defining criteria to identify and remove less crucial parameter groups [32, 16], or a combination of both [23, 56, 13]. The practice of pruning at initialization [14, 30, 7], pruning during training [52, 5, 6, 18], and pruning the pre-trained model [9, 56, 55, 16, 60] also represents an extensively explored area in network pruning.
@@
Regularization in Pruning. Regularization is a common approach to learning sparsity in DNNs [44, 42, 61]. The introduction of a penalty factor controls the level of regularization. When a modest penalty factor is uniformly applied to all weights, it facilitates the gradual acquisition of regular sparsity [57, 13]. Conversely, adopting a larger or progressively increasing penalty factor targeted at a specific parameter group yields large regularization effects [9, 55, 56]. Moreover, while utilizing parameter saliency is a widely adopted method for direct pruning [32, 18], the combination of parameter saliency and regularization is often used for the gradual learning of structural sparsity [9, 55, 56, 13].
@@
One-cycle Pruning. The pruning of a pre-trained network and fine-tuning the resulting pruned architecture is a commonly used framework for network slimming. However, this approach can be computationally intensive due to the multiple rounds of network training it entails. An alternative approach, known as pruning at initialization, quickly gained popularity and includes methods such as Lottery Ticket Hypothesis [14] and its variants [2, 58], single-shot network pruning (SNIP) [30], and progressive skeletonization [7]. However, some limitations of this approach are the risk of sub-optimal network structures, potential performance gaps, and difficulty in defining pruning criteria [14]. Another framework involving pruning and training in just a single training cycle includes methods such as PaT [52], LAASP [18], and only-train-once (OTO) [5, 6]. Although these approaches are more informative and have great potential compared to pruning at initialization, there is little attention in the literature. Therefore, in this paper, we further investigate the one-cycle pruning approach and develop an efficient structured pruning algorithm while achieving high accuracy.
@@ 3.2 Group Saliency Criteria
Conventional saliency estimation for pruning convolutional filters considers only the weight values within this filter to measure its importance. However, in this paper, we also take into account its structurally associated coupled parameters for estimating the saliency, which we call group saliency. The DepGraph [13] algorithm is used to analyze parameter dependencies within $W$, and entire trainable parameters are partitioned into several groups $\mathcal{G}=\{g\}$. Simple norm-based criteria [32] is utilized for group saliency estimation. Each group $g=\{w_{1},w_{2},...,~{}w_{|g|}\}$ has an arbitrary number of sets of parameters with different dimensionality. Given our adoption of global pruning, we propose to incorporate local normalization into the calculation of group importance, thus facilitating the global ranking of filters for the pruning process. The overall saliency score of a group, as determined by the $l_{2}$-norm used in this study, is defined as:
@@
where |.||.|| . | denotes set cardinality.
In Eq. 2, we calculate the normalized $l_{2}$-norm [1] of the set of elements within a group. The resulting local norms are summed up and again normalized with the cardinality of the group. Such normalization is induced because the group size and dimensionality of the set of parameters within the groups could be different. This will ensure fair comparison within different groups or layers during global pruning.
@@ 3.3 Pruning Stability
During the pruning sub-network search process, at the end of each training epoch, we temporarily prune the network for a given pruning constraint and ratio using group saliency criteria defined in Eq. 2. Suppose $\mathcal{M}_{t-i}$ and $\mathcal{M}_{t}$ denotes the pruned sub-network structures at training epoch $t-i$ and $t$, respectively. Those sub-networks are used to check for pruning stability defined by their similarity. As soon as the sub-network stabilizes in subsequent training epochs, the pruning is made permanent, and the resulting pruned sub-network is further trained for convergence.
PaT [52] uses a similar approach, but they check for pruning stability calculated using only the number of channels in each layer. The limitation of their approach is that the two sub-networks could be identical in terms of the number of channels in each layer. However, the identity of those channels derived from the original network could still differ, leading to unstable pruning.
To solve this issue, in this paper, we propose utilizing the total number of retained filters and the identity of derived filters from the original network to check for the pruning stability.
Suppose $\mathcal{F}_{t-i}^{l}$ and $\mathcal{F}_{t}^{l}$ be the set of filters in temporarily pruned sub-networks at training epoch $t-i$ and $t$, respectively. The similarity between these two sub-networks based on the cardinality of the filter set, as well as their corresponding identity, in each layer, is defined using Jaccard Similarity $J$:
@@
The similarity value $J$ ranges from 00 to $1$, where $1$ means two sub-networks are the same and vice-versa. In practice, we use the average of past $r$ similarities:
@@ 3.4 Structured Sparsity Regularization
The structured sparsity regularization scheme in our approach is motivated by pruning methods DepGraph [13] and GReg [56]. DepGraph [13] pruning scheme gradually sparsifies the structurally coupled parameters at the group level based on the normalized group importance scores. During the training process, the regularization is applied to all the parameter groups while driving the groups with low importance scores toward zero for pruning. On the other hand, GReg [56] imposes a penalty term on those filters selected for pruning during the training process while gradually increasing the regularization penalty. However, the GReg [56] technique ignores the structurally coupled parameters while imposing growing regularization for sparsity learning.
@@
[…]
2:  for $t=0$, $\dots$, $T-1$ do
@@
5:     Get pruned model $\mathcal{M}_{t}$ and pruning groups $\mathcal{G}_{\text{prune}}^{t}$ for pruning ratio $\alpha$
@@
[…]
Suppose, at training epoch $t$, $\mathcal{G}_{\text{prune}}^{t}$ and $\mathcal{G}_{\text{non-prune}}^{t}$ denotes the partitioning of structural groups into pruning and non-pruning groups. First, similar to GReg [56], we impose a growing $l_{2}$-norm penalty estimated using a pruning group of parameters to the network’s original loss function defined as:
@@
In the early stage of sparsity learning, Eq. 9 slightly decreases the weight parameters, and this process becomes more aggressive as training progresses. This two-step structured sparsity learning ensures both smooth and fast regularization so that pruning becomes stable in the early stage of network training leaving sufficient time for fine-tuning the pruned model.
@@ 4.1 Experimental Setup
To evaluate our pruning algorithm, we conducted experiments on the CIFAR-10/100 [27] and ImageNet [49] datasets. Initially, we pruned the relatively simple single-branch VGGNet [53] on the CIFAR-10 and CIFAR-100 datasets. Next, we evaluate the pruning of more complex multi-branch ResNet [21] models, on both CIFAR-10 and ImageNet datasets. Finally, we evaluated the performance of pruning a compact network, MobileNetV2 [50], on the ImageNet dataset.
Both CIFAR and ImageNet datasets are trained with a Stochastic Gradient Descend (SGD) optimizer with a momentum of 0.9 using data argumentation strategies adapted from official PyTorch [48] examples. For training on CIFAR-10/100 datasets, the models are trained for 300 epochs with batch size 128, in which the learning rate follows the MultiStepLR scheduler. Conversely, for training on the ImageNet dataset, the learning rate follows a cosine function-shaped decay strategy, gradually approaching zero. The training settings for MobileNetV2 on the ImageNet dataset are adapted from HBONet [31]. Please refer to the supplement for more comprehensive information on the training and pruning parameters employed in our experiments.
For small-scale datasets like CIFAR, slight variations in results across training trials are observed. Thus, we report the average outcome from three trials. Whereas for ImageNet, results are reported from a single training trial.

### v2

# One-Cycle Structured Pruning via Stability-Driven Subnetwork Search

@@ Abstract
Existing structured pruning typically involves multi-stage training procedures that often demand heavy computation. Pruning at initialization, which aims to address this limitation, reduces training costs but struggles with performance. To address these challenges, we propose an efficient one-cycle structured pruning framework that integrates pre-training, pruning, and fine-tuning into a single training cycle without compromising model performance, referred to as the ‘one-cycle approach’. The core idea is to search for the optimal sub-network during the early stages of network training, guided by norm-based group saliency criteria and structured sparsity regularization. We introduce a novel pruning indicator that identifies the stable pruning epoch by measuring the similarity between evolving pruning sub-networks across consecutive training epochs. Additionally, group sparsity regularization helps accelerate the pruning process, thereby speeding up overall training. Extensive experiments on the CIFAR-10/100 and ImageNet datasets using VGGNet, ResNet, and MobileNet architectures demonstrate that our method achieves state-of-the-art accuracy while being one of the most efficient pruning frameworks in terms of training time. Our code is available at https://github.com/ghimiredhikura/OCSPruner.
@@ 1 Introduction
Over the past decade, deep neural networks (DNNs) have demonstrated superior performance at the expense of substantial memory usage, computing power, and energy consumption [he2016deep, deng2020model, zhai2022scaling, ghimire2022survey]. For compressing and deploying these complex models on low-capability devices, pruning has emerged as a particularly promising approach for many embedded and general-purpose computing platforms. Structured pruning [li2017pruning, wang2021neural, chen2021only, Shen_2022_CVPR, fang2023depgraph, chen2023otov2] and unstructured pruning [han2015learning, frankle2018the, ding2019global, sanh2020movement] represent the two predominant methods in pruning. Also, there has been increasing interest in middle-level pruning granularity [lin20221xn, zhang2022learning, park2023balanced], which provides the capability to achieve fine-grained structured sparsity, combining the advantages of both unstructured fine-grained sparsity and structured coarse-grained sparsity.
Figure 1: Pruning ResNet50 on the ImageNet dataset. Each method is shown as a colored circle, with an overlaid star indicating the variant trained from scratch. The marker size reflects the training speedup relative to the baseline training cost.
@@
With the emergence of large-scale DNN models [brown2020language, openai2023gpt4, zhai2022scaling], there is a growing need to reassess the conventional pruning approach of pre-training, pruning, and fine-tuning, which demands excessive resources. Conventional pruning often requires multi-stage pipelines with repeated fine-tuning procedures or offline re-training. Without efficient learning mechanisms, it is not only a waste of computational power and time but also less applicable to real-time scenarios. One-cycle training and pruning, also known as one-cycle pruning, presents itself as a viable alternative solution to this problem. It is therefore important to explore whether a one-cycle pruning framework can quickly adapt to new tasks or environments in real-world applications such as autonomous vehicles, drones, and robots [ral24ghafourian, chen2021only].
To address the limitations of conventional pruning, we introduce One-Cycle Structured Pruner (OCSPruner), which eliminates the requirement for pre-trained base network for pruning. Considering scenarios where networks need to learn from scratch within limited resource budgets, the efficiency gained from avoiding pre-training the base model becomes more compelling. Hence, we focus on efficient and stable structured pruning, ensuring model performance in a more sustainable way. Basically, structured pruning offers the advantage of creating slimmer yet structured networks that do not rely on specific hardware or software; however, it can occasionally lead to performance drops, particularly when dealing with high pruning ratios.
Considering the insights from the Lottery Ticket Hypothesis [frankle2018the], which suggests that sparse, trainable subnets can be identified within larger networks, the timing and method of pruning become critical. Our focus on one-cycle structured pruning involves pruning the network at initialization [frankle2018the, lee2018snip, jorge2021progressive]. However, this type of method may suffer from potential sub-optimal network structures that result in performance degradation. Thus, rather than pruning at initialization, our proposed strategy involves early-phase pruning during training while still maintaining a one-cycle approach. Despite related methods like pruning-aware training (PaT) [Shen_2022_CVPR] and loss-aware automatic selection of pruning criteria (LAASP) [GHIMIRE2023104745], they overlook structural parameter grouping during saliency estimation, leading to reduced performance [fang2023depgraph]. Specifically, PaT [Shen_2022_CVPR] focuses solely on total channel numbers, not considering individual channel identities during stable sub-network selection, while LAASP [GHIMIRE2023104745] lacks stability estimation criteria and involves complex iterative pruning. This paper addresses these limitations and introduces additional algorithmic enhancements for early-phase structured pruning during training from scratch, resulting in state-of-the-art performance.
@@
The main contributions are summarized as follows:
@@
  • A robust pruning paradigm is suggested, which continuously monitors the evolving sub-structure guided by a pruning stability indicator to achieve an optimal architecture during the early stages of network training.
  • The proposed pruning algorithm is validated on VGGNet, ResNet, and MobileNet architectures. Notably, it achieves 75.49% top-1 and 92.63% top-5 accuracy for the ResNet50 model on the ImageNet dataset, with 1.38× faster training and 57% fewer FLOPs.
@@ 2 Related Work
Structured Pruning. While unstructured pruning has been a subject of long-standing research [NIPS1989_6c9882bb, NIPS1992_303ed4c6, han2015learning], structured pruning gained popularity with the emergence of modern DNNs [li2017pruning]. Structured pruning revolves around extracting efficient sub-structures from complete models while optimizing storage and inference processes. This can involve incorporating regularization techniques to induce sparsity [louizos2017learning, ye2018rethinking] or defining criteria to identify and remove less crucial parameter groups [li2017pruning, he2022filter, ghimire2022magnitude], or a combination of both [he2019filter, wang2021neural, fang2023depgraph]. The practice of pruning at initialization [frankle2018the, lee2018snip, jorge2021progressive], pruning during training [Shen_2022_CVPR, chen2021only, chen2023otov2, GHIMIRE2023104745], and pruning the pre-trained model [ding2018auto, wang2021neural, wang2019structured, ghimire2022magnitude, ye2023performance] also represents an extensively explored area in network pruning.
@@
Regularization in Pruning. Regularization is a common approach to learning sparsity in DNNs [louizos2017learning, liu2017learning, ye2018rethinking]. The introduction of a penalty factor controls the level of regularization. When a modest penalty factor is uniformly applied to all weights, it facilitates the gradual acquisition of regular sparsity [wen2016learning, fang2023depgraph]. Conversely, adopting a larger or progressively increasing penalty factor targeted at a specific parameter group yields large regularization effects [ding2018auto, wang2019structured, wang2021neural]. Moreover, while utilizing parameter saliency is a widely adopted method for direct pruning [li2017pruning, GHIMIRE2023104745], the combination of parameter saliency and regularization is often used for the gradual learning of structural sparsity [ding2018auto, wang2019structured, wang2021neural, fang2023depgraph]. While tuning task and model-specific hyperparameters has been challenging in regularization-based pruning, the advent of automatic parameter tuning algorithms using AutoML [falkner2018bohb] and NAS [wang2023prenas] can simplify the process.
@@
One-cycle Pruning. The pruning of a pre-trained network and fine-tuning the resulting pruned architecture is a commonly used framework for network slimming. However, this approach can be computationally intensive due to the multiple rounds of network training it entails. An alternative approach, known as pruning at initialization, quickly gained popularity and includes methods such as Lottery Ticket Hypothesis [frankle2018the] and its variants [bai2022dual, pmlr-v206-xiong23a], single-shot network pruning (SNIP) [lee2018snip], and progressive skeletonization [jorge2021progressive]. However, some of its limitations are the risk of sub-optimal network structures, potential performance gaps, and difficulty in defining pruning criteria [frankle2018the]. Another framework involving pruning and training in just a single training cycle includes methods such as PaT [Shen_2022_CVPR], LAASP [GHIMIRE2023104745], and only-train-once (OTO) [chen2021only, chen2023otov2]. Although these approaches are more informative and have great potential compared to pruning at initialization, there is little attention in the literature. Therefore, in this paper, we further investigate the one-cycle pruning approach and develop an efficient structured pruning algorithm while achieving high accuracy.
@@ 3.2 Group Saliency Criteria
Conventional saliency estimation for pruning convolutional filters considers only the weight values within this filter to measure its importance. However, in this paper, we also take into account its structurally associated coupled parameters for estimating the saliency, which we call group saliency. The DepGraph [fang2023depgraph] algorithm is used to analyze parameter dependencies within $W$, and entire trainable parameters are partitioned into several groups $\mathcal{G}=\{g\}$. Simple norm-based criteria [li2017pruning] is utilized for group saliency estimation. Each group $g=\{w_{1},w_{2},...,~w_{|g|}\}$ has an arbitrary number of sets of parameters with different dimensionality. Given our adoption of global pruning, we propose to incorporate local normalization into the calculation of group importance, thus facilitating the global ranking of filters for the pruning process. The overall saliency score of a group, as determined by the $l_{2}$-norm used in this study, is defined as:
@@
where $|.|$ denotes set cardinality.
In Eq. 2, we calculate the normalized $l_{2}$-norm [NIPS2016_6e7d2da6] of the set of elements within a group. The resulting local norms are summed up and again normalized with the cardinality of the group. Such normalization is induced because the group size and dimensionality of the set of parameters within the groups could be different. This will ensure fair comparison within different groups or layers during global pruning.
@@ 3.3 Pruning Stability
During the pruning sub-network search process, at the end of each training epoch, we temporarily prune the network for a given pruning constraint and ratio using group saliency criteria defined in Eq. 2. Suppose $\mathcal{M}_{t-i}$ and $\mathcal{M}_{t}$ denotes the pruned sub-network structures at training epoch tit-i and $t$, respectively. Those sub-networks are used to check for pruning stability defined by their similarity. As soon as the sub-network stabilizes in subsequent training epochs, the pruning is made permanent, and the resulting pruned sub-network is further trained for convergence.
PaT [Shen_2022_CVPR] uses a similar approach, but they check for pruning stability calculated using only the number of channels in each layer. The limitation of their approach is that the two sub-networks could be identical in terms of the number of channels in each layer. However, the identity of those channels derived from the original network could still differ, leading to unstable pruning. To solve this issue, in this paper, we propose assessing pruning stability using the total number of retained filters and the identity of derived filters from the original network.
Suppose $\mathcal{F}_{t-i}^{l}$ and $\mathcal{F}_{t}^{l}$ be the set of filters in temporarily pruned sub-networks at training epoch tit-i and $t$, respectively. The similarity between these two sub-networks based on the cardinality of the filter set, as well as their corresponding identity, in each layer, is defined using Jaccard Similarity $J$:
@@
The similarity value $J$ ranges from 0 to $1$, where $1$ means two sub-networks are the same and vice-versa. In practice, we use the average of past $r$ similarities:
@@ 3.4 Structured Sparsity Regularization
The structured sparsity regularization scheme in our approach is motivated by pruning methods DepGraph [fang2023depgraph] and GReg [wang2021neural]. DepGraph [fang2023depgraph] pruning scheme gradually sparsifies the structurally coupled parameters at the group level based on the normalized group importance scores. During the training process, the regularization is applied to all the parameter groups while driving the groups with low importance scores toward zero for pruning. On the other hand, GReg [wang2021neural] imposes a penalty term on those filters selected for pruning during the training process while gradually increasing the regularization penalty. However, the GReg [wang2021neural] technique ignores the structurally coupled parameters while imposing growing regularization for sparsity learning.
@@
1:  $\text{SL\_STAGE}\leftarrow$ False
2:  for $t=0$, $\dots$, T1T-1 do
@@
5:   Obtain the pruned model $\mathcal{M}_{t}$ and pruning groups $\mathcal{G}_{\text{prune}}^{t}$ at pruning ratio $\alpha$.
@@
In Eq. 7, $t$ represents the training epoch, and $\delta$ denotes the growing factor. The term $t_{\text{sl-start}}$ corresponds to the starting epoch for structured sparsity regularization. From $t_{\text{sl-start}}$ onward, the penalty factor increases by $\delta$ in every $\Delta t$ epoch interval, as determined by the floor function $\lfloor$$\rfloor$ (Line 17).
Suppose that, at training epoch $t$, $\mathcal{G}_{\text{prune}}^{t}$ and $\mathcal{G}_{\text{non-prune}}^{t}$ denote the partitioning of structural groups into pruning and non-pruning groups. First, similar to GReg [wang2021neural], we impose a growing $l_{2}$-norm penalty estimated using a pruning group of parameters to the network’s original loss function defined as:
@@
In the early stage, sparsity learning, Eq. 9 gently reduces weights, then gradually intensifies. This two-step process provides smooth yet fast regularization, making pruning stable early while leaving enough time for fine-tuning the pruned model.
@@ 4.1 Experimental Setup
To evaluate our pruning algorithm, we conducted experiments on the CIFAR-10/100 [krizhevsky2009learning] and ImageNet [russakovsky2015imagenet] datasets. Initially, we pruned the relatively simple single-branch VGGNet [Simonyan15] on the CIFAR-10 and CIFAR-100 datasets. Next, we evaluate the pruning of more complex multi-branch ResNet [he2016deep] models, on both CIFAR-10 and ImageNet. Finally, we evaluated the pruning of a compact network, MobileNetV2 [sandler2018mobilenetv2], on the ImageNet.
Both CIFAR and ImageNet datasets are trained with a Stochastic Gradient Descend (SGD) optimizer with a momentum of 0.9 using data argumentation strategies adapted from official PyTorch [paszke2017automatic] examples. For training on CIFAR-10/100 datasets, the models are trained for 300 epochs with batch size 128, in which the learning rate follows the MultiStepLR scheduler. Conversely, for training on the ImageNet dataset, the learning rate follows a cosine function-shaped decay strategy, gradually approaching zero. The training settings for MobileNetV2 on the ImageNet dataset are adapted from HBONet [li2019hbonet]. Please refer to the supplement for more comprehensive information on the training and pruning parameters.
For the CIFAR dataset, we observe slight variations across training trials; therefore, we report the average over three runs. In contrast, ImageNet results were consistent across runs; hence, results from a single trial are reported.